# Generalizations of the Squircle-Lemniscate Relation and Keplerian Dynamics

### v2

@@ 2. Sinusoidal Spirals
When $n=1$ the sinusoidal spiral is just the circle with cartesian equation $\left(x-\frac{1}{2}\right)^{2}+y^{2}=\frac{1}{4}$ and when $n=2$, the sinusoidal spiral is just the lemniscate with cartesian equation $(x^{2}+y^{2})^{2}=x^{2}-y^{2}$.
@@ 3. The Lamé-Sinusoidal Spiral Relation
We apply integration by parts to the right-hand side:
$u$ $=(1-x^{2n})^{\frac{1}{2n}}$ $dv$ $=dx$
$du$ $=-x^{2n-1}(1-x^{2n})^{\frac{1}{2n}-1}\,dx$ $v$ $=x$
We obtain
$\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$ $\quad=\quad\left[x(1-x^{2n})^{\frac{1}{2n}}\right]_{0}^{1}-\int_{0}^{1}x\left(-x^{2n-1}(1-x^{2n})^{\frac{1}{2n}-1}\right)\,dx$
$\quad=\quad 0-\int_{0}^{1}\frac{x^{2n}}{(1-x^{2n})^{1-\frac{1}{2n}}}\,dx$
$\quad=\quad\int_{0}^{1}\frac{1-(1-x^{2n})}{(1-x^{2n})^{1-\frac{1}{2n}}}\,dx$
$\quad=\quad\int_{0}^{1}(1-x^{2n})^{\frac{1}{2n}-1}\,dx-\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$
Rearranging terms yields
$2\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$ $\quad=\quad\int_{0}^{1}(1-x^{2n})^{\frac{1}{2n}-1}\,dx$
$\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$ $\quad=\quad\frac{1}{2}\int_{0}^{1}(1-x^{2n})^{\frac{1}{2n}-1}\,dx$
We then make the following substitution in the right hand side of the last equation:
$v=x^{2n}\quad\Longrightarrow\quad dv=2nx^{2n-1}=2nv^{\frac{2n-1}{2n}}dx\quad\Longrightarrow\quad dx=\frac{1}{2n}v^{\frac{1}{2n}-1}dv$
We obtain
$\qquad\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$ $\quad=\quad\frac{1}{2}\int_{0}^{1}\frac{1}{2n}v^{\frac{1}{2n}-1}(1-v)^{\frac{1}{2n}-1}\,dv$
$\quad=\quad\frac{1}{4n}\int_{0}^{1}\left[v(1-v)\right]^{\frac{1}{2n}-1}\,dv$
(10) $\quad=\quad\frac{1}{4n}\left[\int_{0}^{\frac{1}{2}}\left[v(1-v)\right]^{\frac{1}{2n}-1}\,dv+\int_{\frac{1}{2}}^{1}\left[v(1-v)\right]^{\frac{1}{2n}-1}\,dv\right].$
Next we observe that under the substitution w=1vw=1-v, we have
$\int_{\frac{1}{2}}^{1}\left[v(1-v)\right]^{\frac{1}{2n}-1}\,dv$ $\quad=\quad-\int_{\frac{1}{2}}^{0}\left[w(1-w)\right]^{\frac{1}{2n}-1}\,dw$
$\quad=\quad\int_{0}^{\frac{1}{2}}\left[v(1-v)\right]^{\frac{1}{2n}-1}\,dv$
Hence equation (10) becomes
(11) $\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx=\frac{1}{2n}\int_{0}^{\frac{1}{2}}\left[v(1-v)\right]^{\frac{1}{2n}-1}\,dv$
Now make the substitution
$u=4v(1-v)\quad\Longrightarrow\quad du=(4-8v)dv=4(1-2v)dv\quad\Longrightarrow\quad dv=\frac{\,du}{4\sqrt{1-u}},$
and observe that
$1-u=1-4v+4v^{2}=(1-2v)^{2}\quad\Longrightarrow\quad\sqrt{1-u}=1-2v\quad\Longrightarrow\quad dv=\frac{\,du}{4\sqrt{1-u}}.$
Substituting into equation (11) we obtain
$\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$ $\quad=\quad\frac{1}{2n}\int_{0}^{1}\left[\frac{u}{4}\right]^{\frac{1}{2n}-1}\frac{\,du}{4\sqrt{1-u}}$
(12) $\quad=\quad\frac{4^{1-\frac{1}{2n}}}{8n}\int_{0}^{1}\frac{u^{\frac{1}{2n}-1}\,du}{\sqrt{1-u}}$
Finally we make the substitution
$u=r^{2n}\quad\Longrightarrow\quad du=2nr^{2n-1}dr=2nu^{1-\frac{1}{2n}}dr\quad\Longrightarrow\quad u^{\frac{1}{2n}-1}du=2ndr.$
This transforms equation (12) into
(13) $\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$ $\quad=\quad\frac{4^{1-\frac{1}{2n}}}{8n}\int_{0}^{1}\frac{2n\,dr}{\sqrt{1-r^{2n}}}$
$\quad=\quad 4^{-\frac{1}{2n}}\int_{0}^{1}\frac{\,dr}{\sqrt{1-r^{2n}}}$
(14) $\quad=\quad 2^{-\frac{1}{n}}\int_{0}^{1}\frac{\,dr}{\sqrt{1-r^{2n}}}$
This equation rearranges into
$\int_{0}^{1}\frac{\,dr}{\sqrt{1-r^{2n}}}=2^{1/n}\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx,$
which proves the theorem. ∎
@@
The integral on the right hand side of equation (9) is clearly the area of the portion of the Lamé curve $x^{2n}+y^{2n}=1$ contained within the first quadrant. Thus the total area enclosed by the Lamé curve is

### v3

@@ 2. Sinusoidal Spirals
When $n=1$ the sinusoidal spiral is just the circle with cartesian equation $\left(x-\frac{1}{2}\right)^{2}+y^{2}=\frac{1}{4}$ and when $n=2$, the sinusoidal spiral is just the lemniscate with cartesian equation $(x^{2}+y^{2})^{2}=x^{2}-y^{2}$. The rotation $\theta\mapsto\theta-\frac{\pi}{2n}$ converts the polar equation of the sinusoidal spiral to the alternative form $r^{n}=\sin(n\theta)$.
@@ 3. The Lamé-Sinusoidal Spiral Relation
The integral $A=\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx$ represents the area of the portion of the Lamé curve $x^{2n}+y^{2n}=1$ contained within the first quadrant. An alternative way of computing this area is to use the following parametrization of the Lamé curve:
$x=\cos^{\frac{1}{n}}(nt),\ y=\sin^{\frac{1}{n}}(nt),\quad 0\leq t\leq\frac{\pi}{2n}.$
[Note that these are the same as the polar equations of the sinusoidal spiral restricted to the interval $\left[0,\frac{\pi}{2n}\right]$.] Thus we have
$A$ $=\frac{1}{2}\int_{0}^{\frac{\pi}{2n}}\left(x\frac{dy}{dt}-y\frac{dx}{dt}\right)\,dt$
$=\frac{1}{2}\int_{0}^{\frac{\pi}{2n}}\left(\frac{\cos^{\frac{1}{n}}(nt)\sin^{\frac{1}{n}}(nt)\cos(nt)}{\sin(nt)}+\frac{\sin^{\frac{1}{n}}(nt)\cos^{\frac{1}{n}}(nt)\sin(nt)}{\cos(nt)}\right)\,dt$
$=\frac{1}{2}\int_{0}^{\frac{\pi}{2n}}\left(\frac{\cos^{\frac{1}{n}}(nt)\sin^{\frac{1}{n}}(nt)\left(\cos^{2}(nt)+\sin^{2}(nt)\right)}{\cos(nt)\sin(nt)}\right)\,dt$
$=\frac{1}{2^{\frac{1}{n}}}\int_{0}^{\frac{\pi}{2n}}\left(\frac{2^{\frac{1}{n}}\cos^{\frac{1}{n}}(nt)\sin^{\frac{1}{n}}(nt)}{2\cos(nt)\sin(nt)}\right)\,dt$
$=\frac{1}{2^{\frac{1}{n}}}\int_{0}^{\frac{\pi}{2n}}\frac{\sin^{\frac{1}{n}}(2nt)}{\sin(2nt)}\,dt$
Substituting $u=2t$ into the last integral, we obtain
$A$ $=\frac{1}{2^{1+\frac{1}{n}}}\int_{0}^{\frac{\pi}{n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du$
(10) $=\frac{1}{2^{1+\frac{1}{n}}}\left[\int_{0}^{\frac{\pi}{2n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du+\int_{\frac{\pi}{2n}}^{\frac{\pi}{n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du\right]$
Using the substitution $v=\frac{\pi}{n}-u$, we see that
$\int_{\frac{\pi}{2n}}^{\frac{\pi}{n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du$ $=-\int_{\frac{\pi}{2n}}^{0}\frac{\sin^{\frac{1}{n}}(\pi-nv)}{\sin(\pi-nv)}\,dv$
$=\int_{0}^{\frac{\pi}{2n}}\frac{\sin^{\frac{1}{n}}(nv)}{\sin(nv)}\,dv$
$=\int_{0}^{\frac{\pi}{2n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du$
Thus (10) simplifies to
(11) $A=\frac{1}{2^{1+\frac{1}{n}}}\left[2\int_{0}^{\frac{\pi}{2n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du\right]=\frac{1}{2^{\frac{1}{n}}}\int_{0}^{\frac{\pi}{2n}}\frac{\sin^{\frac{1}{n}}(nu)}{\sin(nu)}\,du.$
Making the substitution $r=\sin^{\frac{1}{n}}(nu)$, we obtain
$\frac{dr}{du}=\frac{\sin^{\frac{1}{n}}(nu)\cos(nu)}{\sin(nu)}=\frac{\sin^{\frac{1}{n}}(nu)\sqrt{1-r^{2n}}}{\sin(nu)},$
which implies that
(12) $A=\frac{1}{2^{\frac{1}{n}}}\int_{0}^{1}\frac{\,dr}{\sqrt{1-r^{2n}}}$
Hence
$\int_{0}^{1}\frac{\,dr}{\sqrt{1-r^{2n}}}=2^{\frac{1}{n}}A=2^{\frac{1}{n}}\int_{0}^{1}\sqrt[2n]{1-x^{2n}}\,dx.$
This completes the proof. ∎
@@
As noted in the proof, the integral on the right hand side of equation (9) is the area of the portion of the Lamé curve $x^{2n}+y^{2n}=1$ contained within the first quadrant. Thus the total area enclosed by the Lamé curve is